# Kinematic Tokenization: Optimization-Based Continuous-Time Tokens for Learnable Decision Policies in Noisy Time Series

### v1

@@ 1 Introduction
The adaptation of Foundation Models, which were originally designed for discrete natural language tokens, to continuous time series data presents a fundamental representation challenge. While text is inherently discrete, physical and economic signals are continuous systems viewed through the lens of a discrete sampling clock. Current approaches often rely on patching or direct quantization of raw values. These methods fail to account for measurement uncertainty or the underlying dynamic laws (i.e. the “physics”) governing system evolution between samples.
@@ 1.1 Physics-Informed Tokenization Framework
In this paper, we bridge this gap by situating the time series problem within the framework of Physics-Informed AI. Rather than treating data points as static integers, we treat them as noisy observations of a latent dynamical system subject to stochastic forcing. We introduce a methodology based on Optimization-Based Data Enrichment [4]. As illustrated in Figure 1, our framework fundamentally incorporates physics through two stochastic channels and two coupled manifolds:
@@
A critical distinction of our approach is the nature of the resulting representation. Traditional methods like Kalman filtering are implicit: they produce best estimates only at pre-specified discrete filtering times enabling discrete data correction and higher fidelity (but still discrete) sampling. In contrast, our optimization framework solves for an explicit continuous spline that defines the state $x(t)$ and all its derivatives for all times $t$ in the interval.
@@ 2.1 Foundation Models for Time Series
The success of Large Language Models has spurred the development of Large Observation Models (LOMs). Current State-of-the-Art (SOTA) architectures like PatchTST [7] and iTransformer [6] utilize patching to capture local semantic information. More recent efforts in 2025, such as Google’s TimesFM-ICF [2], have introduced in-context learning to adapt to non-stationary data. However, these models remain physically agnostic; they treat time series patches as statistical vectors, ignoring the underlying kinematic continuity of the signal.
@@ 2.2 Emergent Continuous Architectures (SSMs & LNNs)
To address the computational inefficiency of Transformers on long sequences, State Space Models (SSMs) like Mamba have gained prominence. Recent applications such as MambaStock [9] and TimeMachine [1] leverage Mamba’s selective scan mechanism to learn discretized latent dynamics ($h_{t}=\mathbf{A}h_{t-1}+\mathbf{B}x_{t}$) with linear complexity. Similarly, Liquid Neural Networks [3] and Physics-Informed Koopman Networks [5] explicitly model neuronal activations or latent states via differential equations to handle irregularity.
While effective, these methods typically operate on raw, noisy discrete inputs and must learn the system dynamics (matrix $\mathbf{A}$ or ODE parameters) from scratch. SplineGPT differentiates itself by decoupling the physical extraction from the pattern recognition. We do not ask the model to extract the physics; we explicitly calculate the continuous state (velocity, acceleration, etc) via spline optimization ex-ante, providing the Transformer with pre-computed physically-grounded representations while retaining the training efficiency of standard attention mechanisms.
@@ 2.3 Physics-Informed Finance
The application of Physics-Informed Neural Networks (PINNs) [8] to finance has largely focused on solving known Partial Differential Equations (PDEs). Recent work by Socaciu et al. (2025) [11] demonstrates the utility of PINNs for solving Black-Scholes and Heston equations for option pricing. Deep learning approaches in limit order books typically rely on discrete spatial or temporal structures [10, 12]. However, these approaches are largely deductive—enforcing known laws onto data. Our approach is inductive: we use optimization-based enrichment [4] to derive a latent kinematic state in the asset price itself, enabling alpha generation in trading scenarios where the governing PDE is unknown.
@@ 3.1 Pre-tokenization Transforms
Financial time series, specifically asset prices $P(t)$, are defined on the non-negative real line $\mathbb{R}^{+}$. Direct analysis of raw prices often introduces scaling instabilities. To map the state space to a domain suitable for standard stochastic calculus, we apply a logarithmic transformation $y_{k}=\ln(P(t_{k}))$. This aligns the data with symmetric stochastic processes (e.g., Gaussian process) used in the enrichment phase.
@@ 3.2 Optimization-Based Spline Enrichment
$\mathrm{s.t.}$ $\hskip 7.11317pt\ddot{x}(t)=v(t)$
$\mathrm{s.t.}$ $\hskip 7.11317pt\tilde{y}_{k}=\int_{t_{k-1}}^{t_{k}}x(t)dt+w_{k},\hskip 7.11317ptk=1,...,N.$
We explicitly distinguish the interpretation of (2) from (1). While (1) represents a physical equation of motion, (2) acts as a phenomenological filter. The measurement equation uses log-transformed aggregate volume ($\tilde{y}_{k}=\ln(V_{k})$). Mathematically, treating the integral of the state $x(t)$ as consistent with the log of the sum is a known approximation. However, in the context of deep learning inputs, this heuristic serves a critical engineering purpose: it compresses the dynamic range of volume spikes (which span orders of magnitude) while enforcing temporal continuity via the spline constraints. We find this approximation empirically robust for stabilizing the tokenization of heavy-tailed volume data, preventing outlier gradients from destabilizing the Transformer backbone.
@@
$\ln(P(t))=c_{0,k}+c_{1,k}(t-t_{k})+c_{2,k}\frac{(t-t_{k})^{2}}{2}+c_{3,k}\frac{(t-t_{k})^{3}}{6},$ (3)
where $t\in[t_{k},t_{k+1}]$ and the constants corresponding to this interval are provided by the optimal solution. Similarly, solving (2) generates a piecewise optimal estimate of the Volume Intensity Function $x(t)$ defined by
$x(t)=\tilde{c}_{0,k}+\tilde{c}_{1,k}(t-t_{k})+\tilde{c}_{2,k}\frac{(t-t_{k})^{2}}{2}+\tilde{c}_{3,k}\frac{(t-t_{k})^{3}}{6}+\tilde{c}_{4,k}\frac{(t-t_{k})^{4}}{24},$ (4)
for $t\in[t_{k},t_{k+1}]$. Here, the coefficients $\tilde{c}$ do not represent strict physical derivatives of a fluid, but rather shape parameters of the local volume profile that allow the model to detect demand surges (velocity $\tilde{c}_{1}$) and exhaustion patterns (acceleration $\tilde{c}_{2}$).
@@ 3.4 Stationarity & Normalization
To resolve this, we implement Dual Window Anchoring. For every context window $W$, we subtract the initial price position $c_{0,0}$ and initial volume intensity $\tilde{c}_{0,0}$ from all tokens in the sequence:
$c^{\prime}_{0,k}$ $=c_{0,k}-c_{0,0}$ (7)
$\tilde{c}^{\prime}_{0,k}$ $=\tilde{c}_{0,k}-\tilde{c}_{0,0}$ (8)
@@
Furthermore, the higher-order derivatives exhibit vast scale disparities (e.g., price velocity $c_{1}\approx 10^{-2}$ vs. volume jounce $\tilde{c}_{4}\approx 5.7$). We compute global statistics ($\sigma\in\mathbb{R}^{7}$) for the non-anchored coefficients over the training corpus and apply Z-score normalization to the derivatives ($c_{1},c_{2},c_{3}$) and volume shape parameters ($\tilde{c}_{1},\tilde{c}_{2},\tilde{c}_{3},\tilde{c}_{4}$).
@@ 4.5 Training Objectives & Labeling
To align the model’s kinematic reasoning with actionable trading signals, we employ a ternary classification objective (Buy, Sell, Hold) based on a “Momentum Logic” threshold. We adopt the notation $g_{t}$ for the discrete categorical label to distinguish it from the continuous measurement observations $y_{k}$ used in the enrichment phase.
For a given context window ending at time $t$, the label $g_{t}$ is determined by the return of the target token $c_{0,t+1}$ relative to the last input token $c_{0,t}$:
$g_{t}=\begin{cases}0(\text{Buy})&\text{if }r_{t+1}>\tau\\ 1(\text{Sell})&\text{if }r_{t+1}<-\tau\\ 2(\text{Hold})&\text{otherwise}\end{cases}$ (9)
where $r_{t+1}=c_{0,t+1}-c_{0,t}$ is the log-return over the next interval. In this work, we set the volatility threshold $\tau=0.01$ (1%). This filters out minor noise, forcing the model to only predict “actionable” moves when the kinematic signal suggests a significant price excursion.
@@
$\mathcal{L}=-\sum_{j\in\{0,1,2\}}w_{c}\mathbb{I}(g=j)\log(\hat{g}_{j})$ (10)
Here, $\hat{g}_{j}$ represents the predicted probability assigned by the model to class $j$ (i.e., the output of the final Softmax layer), and $\mathbb{I}(g=j)$ is the indicator function for the true label $g$. We assign weights $\mathbf{w}=[2.0,10.0,1.0]$. The high penalty on the Sell class ($w_{sell}=10.0$) conditions the model to be hyper-sensitive to downside kinematic signatures (e.g., negative jerk), prioritizing capital preservation over accuracy.
@@ 5.2.2 Financial Performance
Having observable behavioral validity, we report the financial outcomes in Tables 5 and 6. SplineGPT reliably outperforms the market benchmark on downside risk (Max Drawdown) while capturing significant upside in momentum regimes (NVDA, TSLA). Notably, in the PFE crash scenario, the model’s ability to recognize negative kinematic signatures (Jerk) allowed it to exit early, resulting in a flat equity curve while the asset declined $\approx 40\%$.
@@
INTC +276.3% 0.0% 0.0% 0.0% +71.6% -37.1% 0.0% 0.0% 0.0% -48.2%
NVDA +544.3% 0.0% 0.0% 0.0% +1185.7% -35.6% 0.0% 0.0% 0.0% -56.4%
JPM +85.8% 0.0% 0.0% 0.0% +159.9% -14.8% 0.0% 0.0% 0.0% -22.1%
XOM +49.2% 0.0% 0.0% 0.0% +28.5% -14.4% 0.0% 0.0% 0.0% -26.3%
PFE +0.4% 0.0% 0.0% 0.0% -41.2% -18.8% 0.0% 0.0% 0.0% -44.5%
TSLA +642.2% 0.0% 0.0% 0.0% +306.2% -33.5% 0.0% 0.0% 0.0% -63.5%
@@
INTC 1.38 N/A N/A N/A 0.85 1.22 N/A N/A N/A 1.12
NVDA 2.28 N/A N/A N/A 2.80 2.11 N/A N/A N/A 3.95
JPM 1.18 N/A N/A N/A 1.60 0.92 N/A N/A N/A 2.25
XOM 0.62 N/A N/A N/A 0.55 0.59 N/A N/A N/A 0.78
PFE -0.26 N/A N/A N/A -0.80 -0.25 N/A N/A N/A -1.05
TSLA 2.15 N/A N/A N/A 1.15 2.13 N/A N/A N/A 1.65
@@ 5.2.4 Kinematic Responsiveness & Regime Adaptation
Figure 6: Cross-Sector Performance Validation (2023 – Present). Left Panel: Equity curves generated by the SplineGPT strategy using Active Risk Management weights. Right Panel: The Buy & Hold benchmark for the same assets. Note the strategy’s ability to smooth out volatility (TSLA) and strictly preserve capital during the PFE crash (red line), yielding a flat return profile while the asset declined 40%. The ablation baselines are omitted because they converged to the Liquidation Equilibrium, realized here by a monotonic Always-Sell action policy; see Tables 5, 6 and 7.
Total Return (Cumulative) Max Drawdown (Risk)
Ticker Spline Patch FD Raw Mkt Spline Patch FD Raw Mkt
INTC +276.3% 0.0% 0.0% 0.0% +71.6% -37.1% 0.0% 0.0% 0.0% -48.2%
NVDA +544.3% 0.0% 0.0% 0.0% +1185.7% -35.6% 0.0% 0.0% 0.0% -56.4%
JPM +85.8% 0.0% 0.0% 0.0% +159.9% -14.8% 0.0% 0.0% 0.0% -22.1%
XOM +49.2% 0.0% 0.0% 0.0% +28.5% -14.4% 0.0% 0.0% 0.0% -26.3%
PFE +0.4% 0.0% 0.0% 0.0% -41.2% -18.8% 0.0% 0.0% 0.0% -44.5%
TSLA +642.2% 0.0% 0.0% 0.0% +306.2% -33.5% 0.0% 0.0% 0.0% -63.5%
Table 5: Absolute Performance Metrics. SplineGPT is compared against (i) an external patching baseline (PatchTST) and (ii) internal discrete ablations (FD, Raw). Across this testbed, all non-spline baselines collapsed to a Liquidation Equilibrium (100% Sell rate), producing a static-cash trajectory (0.0% return, 0.0% drawdown). This confirms that discrete representations fail to generate sufficient signal confidence to overcome the asymmetric risk penalty. SplineGPT reliably escapes this equilibrium, producing an active policy with improved downside control.
Sharpe Ratio Sortino Ratio
Ticker Spline Patch FD Raw Mkt Spline Patch FD Raw Mkt
INTC 1.38 N/A N/A N/A 0.85 1.22 N/A N/A N/A 1.12
NVDA 2.28 N/A N/A N/A 2.80 2.11 N/A N/A N/A 3.95
JPM 1.18 N/A N/A N/A 1.60 0.92 N/A N/A N/A 2.25
XOM 0.62 N/A N/A N/A 0.55 0.59 N/A N/A N/A 0.78
PFE -0.26 N/A N/A N/A -0.80 -0.25 N/A N/A N/A -1.05
TSLA 2.15 N/A N/A N/A 1.15 2.13 N/A N/A N/A 1.65
Table 6: Risk-Adjusted Efficiency. Risk-adjusted metrics are undefined for baselines that take no positions (zero variance returns). Under the asymmetric action objective, PatchTST, FD, and Raw converge to an abstention equilibrium and therefore yield N/A Sharpe/Sortino. SplineGPT yields non-trivial action and measurable efficiency, supporting the claim that the continuous, coupled kinematic representation improves the learnability of selective entry/exit behavior in this setting.

### v2

@@ 1 Introduction
The adaptation of Foundation Models, which were originally designed for discrete natural language tokens, to continuous time series data presents a fundamental representation challenge. While text is inherently discrete, physical and economic signals are continuous systems viewed through the lens of a discrete sampling clock. Current approaches often rely on patching or direct quantization of raw values. These methods fail to account for measurement uncertainty or the underlying dynamic laws (i.e. the “physics”) governing system evolution between samples. While financial markets inherently lack the rigorous first-principle governance found in natural physical systems, in our experiments we impose Newton-like equations of motion on the evolution of trading parameters, utilizing kinematic consistency as a robust inductive bias for representation learning.
@@ 1.1 Physics-Informed Tokenization Framework
In this paper, we bridge this gap between discrete measurement samples and the continuous underlying processes by situating the time series problem within a framework of Physics-Informed AI. Rather than treating data points as static numbers, we treat them as noisy observations of a latent dynamical system subject to stochastic forcing. We introduce a methodology based on Optimization-Based Data Enrichment [4]. As illustrated in Figure 1, our framework fundamentally incorporates physics through two stochastic channels and two coupled manifolds:
@@
A critical distinction of our approach is the nature of the resulting representation. Traditional methods like Kalman filtering are implicit: they produce best estimates at pre-specified discrete filtering times enabling discrete data correction and higher fidelity (but still discrete) sampling. Continuous limits of the Kalman approach do exist [11], but these treatments fail to produce compact discrete representations (e.g. spline coefficients) which are natural for use in Transformer architectures; these methods are typically implicit and recursive. In contrast, our optimization framework solves for an explicit continuous spline that defines the state $x(t)$ and all its derivatives for all times $t$ in the interval via a direct compact representation.
@@ 2.1 Foundation Models for Time Series
The success of Large Language Models has spurred the development of Large Observation Models (LOMs). Current state-of-the-art architectures like PatchTST [8] and iTransformer [7] utilize patching to capture local semantic information. More recent efforts in 2025, such as Google’s TimesFM-ICF [2], have introduced in-context learning to adapt to non-stationary data. However, these models remain physically agnostic; they treat time series patches as statistical vectors, ignoring the underlying kinematic continuity of the signal.
@@ 2.2 Emergent Continuous Architectures (SSMs & LNNs)
To address the computational inefficiency of Transformers on long sequences, State Space Models (SSMs) like Mamba have gained prominence. Recent applications such as MambaStock [12] and TimeMachine [1] leverage Mamba’s selective scan mechanism to learn discretized latent dynamics with linear complexity. Similarly, Liquid Neural Networks [3] and Physics-Informed Koopman Networks [6] explicitly model neuronal activations or latent states via differential equations to handle irregularity.
While effective, these methods typically operate on raw, noisy discrete inputs and must learn the system dynamics (matrix $\mathbf{A}$ or ODE parameters) from scratch. SplineGPT differentiates itself by decoupling the physical extraction from the pattern recognition. We do not ask the model to extract the physics; we explicitly calculate a continuous state (velocity, acceleration, etc) via spline optimization ex-ante, providing the Transformer with pre-computed physically-grounded discrete representations of the continuous physics while retaining the training efficiency of standard attention mechanisms.
@@ 2.3 Physics-Informed Finance
The application of Physics-Informed Neural Networks (PINNs) [9] to finance has largely focused on solving known Partial Differential Equations (PDEs). Recent work by Socaciu et al. (2025) [14] demonstrates the utility of PINNs for solving Black-Scholes and Heston equations for option pricing. Deep learning approaches in limit order books typically rely on discrete spatial or temporal structures [13, 15]. However, these approaches are largely deductive—enforcing known laws onto data. Our approach is inductive: we use optimization-based enrichment [4] to derive a latent kinematic state from the asset price data stream, enabling alpha generation in trading scenarios where a general governing law is unknown.
It should be noted that in the present approach local “laws” are prescribed in the absence of fundamental physics (e.g. stochastically forced second order motion mirroring Newton’s second law) which are used for the financial data tokenization. We emphasize that the tokenization scheme that is described here is adaptable to other data domains where rich physical frameworks governing system behavior are proven by modifying the governing laws and measurement models. The same general treatment is applicable leveraging rigorous physics when available to adjust the equivalent data enrichment formulation as we demonstrate in (1) and (2).
@@ 3.1 Pre-tokenization Transforms
Financial time series, specifically asset prices $P(t)$, are defined on the non-negative real line $\mathbb{R}^{+}$. Direct analysis of raw prices often introduces scaling instabilities. To map the state space to a domain better suited to standard stochastic processes, we apply a logarithmic transformation $y_{k}=\ln(P(t_{k}))$. This aligns the data with symmetric stochastic processes (e.g., Gaussian process) used in the enrichment phase.
@@ 3.2 Optimization-Based Spline Enrichment
$\underset{\tilde{v},\tilde{w},\tilde{x}}{\min}$ $\hskip 7.11317pt\int_{t_{0}}^{t_{N}}\frac{1}{2}\tilde{v}^{2}dt+\frac{\alpha^{2}}{2}\sum_{k}\tilde{w}_{k}^{2}$ (2)
$\mathrm{s.t.}$ $\hskip 7.11317pt\ddot{\tilde{x}}(t)=\tilde{v}(t)$
$\mathrm{s.t.}$ $\hskip 7.11317pt\tilde{y}_{k}=\int_{t_{k-1}}^{t_{k}}\tilde{x}(t)dt+\tilde{w}_{k},\hskip 7.11317ptk=1,...,N.$
We explicitly distinguish the interpretation of (2) from (1). While (1) represents a physical equation of motion, (2) acts as a phenomenological filter [10]. The measurement equation uses log-transformed aggregate volume ($\tilde{y}_{k}=\ln(V_{k})$). Mathematically, treating the integral of the state $\tilde{x}(t)$ as consistent with the log of the sum is a known approximation. However, in the context of deep learning inputs, this heuristic serves a critical engineering purpose: it compresses the dynamic range of volume spikes (which span orders of magnitude) while enforcing temporal continuity via the spline constraints. We find this approximation empirically robust for stabilizing the tokenization of heavy-tailed volume data, preventing outlier gradients from destabilizing the Transformer backbone.
@@
$x(t)=c_{0}(t_{k})+c_{1}(t_{k})(t-t_{k})+c_{2}(t_{k})\frac{(t-t_{k})^{2}}{2}+c_{3}(t_{k})\frac{(t-t_{k})^{3}}{6},$ (3)
where $t\in[t_{k},t_{k+1}]$ and the constants corresponding to this interval are provided by the optimal solution. The coefficients $c_{j}(t_{k})$ represents the $j$-th kinematic term (e.g. position, velocity, etc) on the interval starting at $t_{k}$. Similarly, solving (2) generates a piecewise optimal estimate of the Volume Intensity Function $\tilde{x}(t)$ defined by
$\tilde{x}(t)=\tilde{c}_{0}(t_{k})+\tilde{c}_{1}(t_{k})(t-t_{k})+\tilde{c}_{2}(t_{k})\frac{(t-t_{k})^{2}}{2}+\tilde{c}_{3}(t_{k})\frac{(t-t_{k})^{3}}{6}+\tilde{c}_{4}(t_{k})\frac{(t-t_{k})^{4}}{24},$ (4)
for $t\in[t_{k},t_{k+1}]$. Here, the coefficients $\tilde{c}$ do not represent strict physical derivatives, but rather shape parameters of the local log-volume rate profile that allow the model to detect demand surges (velocity $\tilde{c}_{1}$) and exhaustion patterns (acceleration $\tilde{c}_{2}$).
@@ 3.4 Stationarity & Normalization
To resolve this, we implement Dual Window Anchoring. For every context window $W$, we subtract the initial price position $c_{0}(t_{0})$ and initial volume intensity $\tilde{c}_{0}(t_{0})$ from all tokens in the sequence:
$c^{\prime}_{0}(t_{k})$ $=c_{0}(t_{k})-c_{0}(t_{0})$ (7)
$\tilde{c}^{\prime}_{0}(t_{k})$ $=\tilde{c}_{0}(t_{k})-\tilde{c}_{0}(t_{0})$ (8)
@@
Furthermore, the higher-order derivatives exhibit vast scale disparities (e.g., price velocity $c_{1}\approx 10^{-2}$ vs. volume jounce $\tilde{c}_{4}\approx 5.7$). We compute global statistics ($\sigma\in\mathbb{R}^{7}$) for the non-anchored coefficients over the training corpus and apply Z-score normalization [5] to the derivatives ($c_{1},c_{2},c_{3}$) and volume shape parameters ($\tilde{c}_{1},\tilde{c}_{2},\tilde{c}_{3},\tilde{c}_{4}$).
@@ 4.5 Training Objectives & Labeling
To align the model’s kinematic reasoning with actionable trading signals, we employ a ternary classification objective (Buy, Sell, Hold) based on a “Momentum Logic” threshold. We adopt the notation $g_{j}$ for the discrete categorical label to distinguish it from the measurement observations $y_{k}$ used in the enrichment phase.
For a given context window ending at time $t_{j}$, the label $g_{j}$ is determined by the state of the target token $c_{0}(t_{j+1})$ relative to the last input token $c_{0}(t_{j})$:
$g_{j}=\begin{cases}0(\text{Buy})&\text{if }r_{j}>\tau\\ 1(\text{Sell})&\text{if }r_{j}<-\tau\\ 2(\text{Hold})&\text{otherwise}\end{cases}$ (9)
where $r_{j}=c_{0}(t_{j+1})-c_{0}(t_{j})$ is the log-return over the next interval. In this work, we set the volatility threshold $\tau=0.01$ (1%). This filters out minor noise, forcing the model to only predict “actionable” moves when the kinematic signal suggests a significant price excursion.
@@
$\mathcal{L}=-\sum_{i\in\{0,1,2\}}w_{i}\mathbb{I}(g=i)\log(\hat{g}_{i})$ (10)
Here, $\hat{g}_{i}$ represents the predicted probability assigned by the model to class $i$, and $\mathbb{I}(g=i)$ is the indicator function for the true label $g$. We assign weights $\mathbf{w}=[2.0,10.0,1.0]$. The high penalty on the Sell class ($w_{sell}=10.0$) conditions the model to be hyper-sensitive to downside kinematic signatures (e.g., negative jerk), prioritizing capital preservation over accuracy.
@@ 5.2.2 Financial Performance
Having established behavioral validity, we report the financial outcomes in Tables 3 and 4, with cumulative equity curves visualized in Figure 6. SplineGPT reliably outperforms the market benchmark on downside risk (Max Drawdown) while capturing significant upside in momentum regimes (NVDA, TSLA). Notably, in the PFE crash scenario, the model’s ability to recognize negative kinematic signatures (Jerk) allowed it to exit early, resulting in a flat equity curve while the asset declined $\approx 40\%$.
@@
INTC +240.4% 0.0% 0.0% 0.0% +71.6% -34.7% 0.0% 0.0% 0.0% -48.2%
NVDA +579.2% 0.0% 0.0% 0.0% +1185.7% -32.1% 0.0% 0.0% 0.0% -56.4%
JPM +82.0% 0.0% 0.0% 0.0% +159.9% -12.5% 0.0% 0.0% 0.0% -22.1%
XOM +48.4% 0.0% 0.0% 0.0% +28.5% -13.5% 0.0% 0.0% 0.0% -26.3%
PFE +3.6% 0.0% 0.0% 0.0% -41.2% -18.8% 0.0% 0.0% 0.0% -44.5%
TSLA +699.8% 0.0% 0.0% 0.0% +306.2% -33.4% 0.0% 0.0% 0.0% -63.5%
@@
INTC 1.24 N/A N/A N/A 0.85 1.13 N/A N/A N/A 1.12
NVDA 2.31 N/A N/A N/A 2.80 2.16 N/A N/A N/A 3.95
JPM 0.98 N/A N/A N/A 1.60 0.88 N/A N/A N/A 2.25
XOM 0.61 N/A N/A N/A 0.55 0.61 N/A N/A N/A 0.78
PFE -0.19 N/A N/A N/A -0.80 -0.18 N/A N/A N/A -1.05
TSLA 2.20 N/A N/A N/A 1.15 2.23 N/A N/A N/A 1.65
@@ 5.2.4 Kinematic Responsiveness & Regime Adaptation
Figure 6: Cross-Sector Performance Validation (2023 – Present). Left Panel: Equity curves generated by the SplineGPT strategy using Active Risk Management weights. Right Panel: The Buy & Hold benchmark for the same assets. Note the strategy’s ability to smooth out volatility (TSLA) and strictly preserve capital during the PFE crash (red line), yielding a flat return profile while the asset declined 40%. The ablation baselines are omitted because they converged to the Liquidation Equilibrium, realized here by a monotonic Always-Sell action policy; see Tables 3, 4 and 5.